A shaft is turning at 65 rad/s at time t=0. Thereafter, its angular acceleration is given by
α = -10 rad/s2 - 5t rad/s3 ,
where t is the elapsed time. (a) Find its angular speed at t=3 s. (b) How far does it turn in these 3s?
Answer #1

The initial angular speed of shaft 1s, a- 65 rad/s a) Angular acceleration is, α=-10-4.5t 10-5t dt 2 At t=0, ω=65 rad/s 65-0+b) Angular displacement is, dt -10--2.5-+65t 0 =1-5(3)-0.83(3) +65(3) = 122 rad Hence, the angular displacement is 122 rad.